what is the mass of impurity in a sample of CaCO3 if 0.040mole of CO2 was formed when 5.0g of the sample reacted with excessbdilute HCL


1
Expert's answer
2021-02-22T06:08:36-0500

CaCO3 + 2HCl → CaCl2 + CO2 + H2O

n(CO2) = n(CaCO3) = 0.04 mol

M(CaCO3) = 100.08 g/mol

"m = n \times M"

m(CaCO3) "= 0.04 \times 100.08 = 4.00 \;g"

Δm = 5.0 – 4.0 = 1.0 g

Answer: 1.0 g